In 1920 the average life expectancy for a man was 53.6 years. In 1965, the average life expectancy for a man was 66.8 years.
a) What two points can you use to compare years to life expectancy? (I did this and got (1920, 53.6) and (1965, 66.8) is that right?)
b) Find and equation for the information (I did this and got 66.8-53.6=13.2 and 1965-1920=45 then got 13.2/45 is that right?)
c) what is the average life expectancy of a man in 2015?
Thank you!!

Answer by stanbon(75887)

In 1920 the average life expectancy for a man was 53.6 years. In 1965, the average life expectancy for a man was 66.8 years.
a) What two points can you use to compare years to life expectancy? (I did this and got (1920, 53.6) and (1965, 66.8) is that right?)
b) Find and equation for the information (I did this and got 66.8-53.6=13.2 and 1965-1920=45 then got 13.2/45 is that right?)
Ans: That is the slope of the equation.
You need the intercept::
Form: y = mx+b
Solve for "b":
53.6 = (13.2/45)*20 + b
b = 53.6 - (13.2/45)20 = 53.59
---
Equation: f(x) = (13.2/45)x+53.59
---------------------------------
c) what is the average life expectancy of a man in 2015?
Let x = 15-20 = -5
f(-5) = (13.2/45)(-5)+53.59 = 52.12 years
=========================
Cheers,
Stan H.
--------------------------